OpenStudy (anonymous):

7 - (2x + 11) + 3(3 - x) = 20 Solve for x

OpenStudy (anonymous):

PEMDAS First, get rid of the parenthesis by distibuting (factoring) outside coeffecient into each term

OpenStudy (anonymous):

7 - 2x - 11 + 9 - 3x = 20 Now combine like terms and isolate X on one side

OpenStudy (anonymous):

Is the answer -3?

OpenStudy (anonymous):

5 - 5x = 20 -15 = 5x x = -3 yes that is correct